# PointAD+: Learning Hierarchical Representations for ZS 3D Anomaly Detection

### v3

@@ 2 Related Work
3D anomaly detection has recently gained attention in the community of anomaly detection. Unlike 2D anomaly detection, which leverages RGB information, 3D anomaly detection focuses on identifying unusual spatial relationships in 3D objects. While 2D anomaly detection has been extensively explored, research on 3D anomaly detection remains relatively limited. This disparity is largely due to the challenges associated with acquiring high-quality 3D data, as collecting 3D datasets is more costly and labor-intensive compared to 2D images. MVTec 3D-AD [20] is the first public dataset, which provides the point cloud anomalies and the corresponding 2D-RGB information. Subsequently, Eyecandies [39] and Real3D-AD [40] provide the 3D dataset for a wide exploration [41]. These datasets bridge the connection between 3D and 2D anomaly detection. 3D-ST [20] uses a teacher net to extract dense local geometry descriptors and design a student net to match such descriptors. AST [21] introduces asymmetric teacher and student net to further improve 3D anomaly detection. Instead of only using point clouds, BTF [19], multimodal 3DM [23, wang2025m3dm], CPFM [43], and SDM [25] integrate point features and RGB pixel features to detect 3D anomalies. While these approaches exhibit commendable performance by storing object-specific normal point and pixel features within the unsupervised learning framework, such paradigms simultaneously limit their generalization capacity to point clouds from unseen objects, which is crucial to detecting anomalies when the target object is unavailable. To fill this gap, we introduce PointAD+ and PointAD to identify unseen anomalies across diverse objects. They extend CLIP to the realm of ZS 3D and show robust generalization in capturing generic normality and abnormality within points. Furthermore, PointAD+ and PointAD serve as a unified framework, allowing seamless integration of 3D and RGB without additional training.
@@ 3.2 Overview of PointAD and PointAD+
We then propose hierarchical representation learning (Section 3.6), which aligns these representations with learnable hierarchical text prompts, enabling the model to incorporate generic anomaly semantics from rendering-based abnormality (rendering/bottom layer) and geometry-based abnormality (geometry /top layer). On the rendering layer (Section 3.6.1), we introduce hybrid representation learning to align rendering prompts with both pixel representations from 2D renderings and implicit point representations in the implicit point space. On the geometry layer (Section 3.6.2), we align geometric prompts with explicit point representations, which are derived from implicit representations and further are encoded in geometry information, in explicit point space.
@@ 3.6 Hierarchical representation learning
The key to ZS 3D anomaly detection lies in the model’s ability to capture generic anomaly semantics, enabling it to detect anomalies across various objects. The above section focuses on effectively representing 3D anomalies. Here, we propose hierarchical representation learning to align these representations on two layers: the rendering/bottom layer aligns the implicit point representation and pixel representation with the learnable rendering prompt in implicit point space; the geometric/top layer aligns the explicit point representations with the learnable geometric prompts in explicit point space. Furthermore, we introduce cross-hierarchy alignment to facilitate mutual anomaly learning across layers. This enables PointAD+ to produce generic text embeddings for generalized anomaly semantics.
@@ 3.6.1 Implicit point representation learning on the rendering layer (PointAD)
$L_{I(3d)}^{g}=\frac{1}{N}\sum_{i}\textrm{CE}(\frac{1}{K}\textstyle\sum_{f_{i}^{(k)}\in p_{i}}P(g^{r}_{c},f_{i}^{(k)}),\max{(y^{3d}_{i})}).$ (6)
@@
$\hskip-2.84544pt\textstyle L_{I(3d)}^{l}=\frac{1}{N}\sum_{i}\Big(\textrm{Dice}(S_{i(n)}^{3d},I-y_{i}^{3d})+\textrm{Dice}(S_{i(a)}^{3d},y_{i}^{3d})\Big).$ (11)
MTL-based pixel representation learning. We further improve PointAD+ and PointAD point understanding by capturing 2D glocal anomaly semantics into the object-agnostic text prompt template. We treat the anomaly recognition for one rendering from the point cloud as a task. Hence, we formulate the anomaly semantics learning for multiple 2D renderings as MTL. MTL-based 2D representation learning is divided into two parts for respective alignment to 2D global and local anomaly semantics. For 2D global semantics, we use CrossEntropy to quantify the discrepancy between the textual representations and each global Pixel representation $f_{i}$. Global MTL-based 2D representation learning $L_{I(2d)}^{g}$ is defined as:
$\textstyle L_{I(2d)}^{g}=\frac{1}{NK}\sum_{i,k}\textrm{CE}(P(g^{r}_{c},f_{i}^{(k)}),\max{(y_{i}^{(k)})}).$ (12)
Also, we focus on 2D abnormal regions to understand pixel-level anomalies. As the anomaly regions are typically smaller than normal regions, we employ Focal Loss to mitigate the class imbalance besides Dice Loss. Let $\oplus$ denote the concatenation operation. Local MTL-based 2D representation learning $L_{I(2d)}^{l}$ is given as follows:
$\begin{split}\textstyle L_{I(2d)}^{l}=\frac{1}{NK}\sum_{i,k}\textrm{Focal}(S_{i(n)}^{2d(k)}\oplus S_{i(a)}^{2d(k)},y_{i}^{(k)})+\\ \textrm{Dice}(S_{i(n)}^{2d(k)},I-y_{i}^{(k)})+\textrm{Dice}(S_{i(a)}^{2d(k)},y_{i}^{(k)}).\end{split}$ (13)
@@
${\color[rgb]{0,0,0}L^{r}_{hybrid}=L_{I(3d)}^{g}+L_{I(3d)}^{l}+L_{I(2d)}^{g}+L_{I(2d)}^{l}.}$ (14)
@@ 3.6.2 Explicit point representation learning on the geometry layer
Here, to incorporate the abnormal spatial relations within point clouds, we minimize $L_{E-3d}^{local}$ ($L^{g}$) to align explicit local point representation $q_{i}^{m}$ with the 3D anomaly semantics:
$\color[rgb]{0,0,0}\textstyle L_{E-3d}^{local}=\frac{1}{N}\sum_{i}\Big(\textrm{Dice}(T_{i(n)}^{3d},I-y_{i}^{3d})+\textrm{Dice}(T_{i(a)}^{3d},y_{i}^{3d})\Big).$} (17)
@@ 4.3.5 Complexity analysis
The computation overhead is also an indicator to assess the effectiveness of the proposed method. We compare the computation overhead in Table VI, including GFLOPs, frames per second (FPS), and GPU memory consumption with a batch size of 1. For a fair comparison, we maintain NVIDIA RTX 3090 24GB GPU idle. PointAD+ requires only about 2.0 more GFLOPs than PointAD (1658.34 vs. 1656.34) yet achieves a significant performance improvement. In contrast, other CLIP-based methods such as AnomalyCLIP and PointCLIP V2 consume comparable or even fewer GFLOPs but deliver worse results. Although CLIP+Rendering and Cheraghian use fewer GFLOPs, they fail to provide competitive performance. The performance and computational overhead of PointAD and PointAD+ are comparable to those of unsupervised methods. As for ZS detection, compared to PointCLIP V2, PointAD requires less time to infer an image, achieving a higher 2.52 vs. 0.66 FPS with lower graphic memory usage. PointAD+ needs a slightly higher computation overhead over PointAD, i.e., 2.52 vs. 2.18 FPS, and attains a noticeable performance improvement on ZS 3D and multimodal 3D. CLIP + rendering has the fastest inference speed, but its detection performance lags behind PointAD+ and PointAD. Therefore, PointAD+ achieves a favorable trade-off between performance and overhead.
@@
$L_{I(3d)}^{g}$ $L_{I(3d)}^{l}$ $L_{I(2d)}^{g}$ $L_{I(2d)}^{l}$ F S Local Global Local Global
@@
(a) The effect of $\alpha$ to the overall performance.
@@ 5.1 Ablation on modules and key hyperparameters
Ablation on $\alpha$. The coefficient $\alpha$ balances the feature fusion between the F and S stages. We investigate the effect of the coefficient $\alpha$ on the model’s performance. We set $\alpha\in\{0,0.1,0.3,0.5,0.7,0.9,1\}$. As shown in Figure 10(a), increasing $\alpha$ initially improves both image-/pixel- auroc performance. However, as $\alpha$ continues to increase, performance begins to decline. Empirically, $\alpha=0.5$ achieves a good balance between image-level and pixel-level performance.

### v4

@@ 2 Related Work
3D anomaly detection has recently gained attention in the community of anomaly detection. Unlike 2D anomaly detection, which leverages RGB information, 3D anomaly detection focuses on identifying unusual spatial relationships in 3D objects. While 2D anomaly detection has been extensively explored, research on 3D anomaly detection remains relatively limited. This disparity is largely due to the challenges associated with acquiring high-quality 3D data, as collecting 3D datasets is more costly and labor-intensive compared to 2D images. MVTec 3D-AD [20] is the first public dataset, which provides the point cloud anomalies and the corresponding 2D-RGB information. Subsequently, Eyecandies [39] and Real3D-AD [40] provide the 3D dataset for a wide exploration [41]. These datasets bridge the connection between 3D and 2D anomaly detection. 3D-ST [20] uses a teacher net to extract dense local geometry descriptors and design a student net to match such descriptors. AST [21] introduces asymmetric teacher and student net to further improve 3D anomaly detection. Instead of only using point clouds, BTF [19], multimodal 3DM [23, 42], CPFM [43], and SDM [25] integrate point features and RGB pixel features to detect 3D anomalies. While these approaches exhibit commendable performance by storing object-specific normal point and pixel features within the unsupervised learning framework, such paradigms simultaneously limit their generalization capacity to point clouds from unseen objects, which is crucial to detecting anomalies when the target object is unavailable. To fill this gap, we introduce PointAD+ and PointAD to identify unseen anomalies across diverse objects. They extend CLIP to the realm of ZS 3D and show robust generalization in capturing generic normality and abnormality within points. Furthermore, PointAD+ and PointAD serve as a unified framework, allowing seamless integration of 3D and RGB without additional training.
@@ 3.2 Overview of PointAD and PointAD+
We then propose hierarchical representation learning (Section 3.6), which aligns these representations with learnable hierarchical text prompts, enabling the model to incorporate generic anomaly semantics from rendering-based abnormality (rendering (bottom) layer) and geometry-based abnormality (geometry (top) layer). On the rendering layer (Section 3.6.1), we introduce hybrid representation learning to align rendering prompts with both pixel representations from 2D renderings and implicit point representations in the implicit point space. On the geometry layer (Section 3.6.2), we align geometric prompts with explicit point representations, which are derived from implicit representations and further are encoded in geometry information, in explicit point space.
@@ 3.6 Hierarchical representation learning
The key to ZS 3D anomaly detection lies in the model’s ability to capture generic anomaly semantics, enabling it to detect anomalies across various objects. The above section focuses on effectively representing 3D anomalies. Here, we propose hierarchical representation learning to align these representations on two layers: the rendering (bottom) layer aligns the implicit point representation and pixel representation with the learnable rendering prompt in implicit point space; the geometric (top) layer aligns the explicit point representations with the learnable geometric prompts in explicit point space. Furthermore, we introduce cross-hierarchy alignment to facilitate mutual anomaly learning across layers. This enables PointAD+ to produce generic text embeddings for generalized anomaly semantics.
@@ 3.6.1 Implicit point representation learning on the rendering layer (PointAD)
$\hskip-2.84544pt\textstyle L_{I(3d)}^{global}=\frac{1}{N}\sum_{i}\textrm{CE}(\frac{1}{K}\textstyle\sum_{f_{i}^{(k)}\in p_{i}}P(g^{r}_{c},f_{i}^{(k)}),\max{(y^{3d}_{i})}).$ (6)
@@
$\hskip-2.84544pt\textstyle L_{I(3d)}^{local}=\frac{1}{N}\sum_{i}\Big(\textrm{Dice}(S_{i(n)}^{3d},I-y_{i}^{3d})+\textrm{Dice}(S_{i(a)}^{3d},y_{i}^{3d})\Big).$ (11)
MTL-based pixel representation learning. We further improve PointAD+ and PointAD point understanding by capturing 2D glocal anomaly semantics into the object-agnostic text prompt template. We treat the anomaly recognition for one rendering from the point cloud as a task. Hence, we formulate the anomaly semantics learning for multiple 2D renderings as MTL. MTL-based 2D representation learning is divided into two parts for respective alignment to 2D global and local anomaly semantics. For 2D global semantics, we use CrossEntropy to quantify the discrepancy between the textual representations and each global Pixel representation $f_{i}$. Global MTL-based 2D representation learning $L_{I(2d)}^{global}$ is defined as:
$\textstyle L_{I(2d)}^{global}=\frac{1}{NK}\sum_{i,k}\textrm{CE}(P(g^{r}_{c},f_{i}^{(k)}),\max{(y_{i}^{(k)})}).$ (12)
Also, we focus on 2D abnormal regions to understand pixel-level anomalies. As the anomaly regions are typically smaller than normal regions, we employ Focal Loss to mitigate the class imbalance besides Dice Loss. Let $\oplus$ denote the concatenation operation. Local MTL-based 2D representation learning $L_{I(2d)}^{local}$ is given as follows:
$\begin{split}\textstyle L_{I(2d)}^{local}=\frac{1}{NK}\sum_{i,k}\textrm{Focal}(S_{i(n)}^{2d(k)}\oplus S_{i(a)}^{2d(k)},y_{i}^{(k)})+\\ \textrm{Dice}(S_{i(n)}^{2d(k)},I-y_{i}^{(k)})+\textrm{Dice}(S_{i(a)}^{2d(k)},y_{i}^{(k)}).\end{split}$ (13)
@@
${\color[rgb]{0,0,0}L^{r}_{hybrid}=L_{I(3d)}^{global}+L_{I(3d)}^{local}+L_{I(2d)}^{global}+L_{I(2d)}^{local}.}$ (14)
@@ 3.6.2 Explicit point representation learning on the geometry layer
Here, to incorporate the abnormal spatial relations within point clouds, we minimize $L^{geo}$ to align explicit local point representation $q_{i}^{m}$ with the 3D anomaly semantics:
$\color[rgb]{0,0,0}\hskip-5.69046pt\textstyle L^{geo}=\frac{1}{N}\sum_{i}\Big(\textrm{Dice}(T_{i(n)}^{3d},I-y_{i}^{3d})+\textrm{Dice}(T_{i(a)}^{3d},y_{i}^{3d})\Big).$} (17)
@@ 4.3.5 Complexity analysis
The computation overhead is also an indicator to assess the effectiveness of the proposed method. We compare the computation overhead in Table VI, including GFLOPs, frames per second (FPS), and GPU memory consumption with a batch size of 1. For a fair comparison, we maintain NVIDIA RTX 3090 24GB GPU idle. PointAD+ requires only about 2.0 more GFLOPs than PointAD (1658.34 vs. 1656.34) yet achieves a significant performance improvement. In contrast, other CLIP-based methods, such as PointCLIP V2, consume larger GFLOPs but deliver worse results. Although CLIP+Rendering, Cheraghian, and AnomalyCLIP use fewer GFLOPs, they fail to provide competitive performance. As for inference speed, compared to PointCLIP V2, PointAD requires less time to infer an image and achieves a higher 2.52 vs. 0.66 FPS with lower graphic memory usage. PointAD+ needs a slightly higher computation overhead over PointAD, i.e., 2.52 vs. 2.18 FPS, and attains a noticeable performance improvement on ZS 3D and multimodal 3D. CLIP + rendering has the fastest inference speed, but its detection performance lags behind PointAD+ and PointAD. From the above results, PointAD+ achieves a favorable trade-off between performance and overhead.
@@
$L_{I(3d)}^{global}$ $L_{I(3d)}^{local}$ $L_{I(2d)}^{global}$ $L_{I(2d)}^{local}$ F S Local Global Local Global
@@
(a) The effect of $\alpha$ to the segmentation performance.
@@ 5.1 Ablation on modules and key hyperparameters
Ablation on $\alpha$. The coefficient $\alpha$ balances the feature fusion between the F and S stages. We investigate the effect of the coefficient $\alpha$ on the model’s performance. We set $\alpha\in\{0,0.1,0.3,0.5,0.7,0.9,1\}$. As shown in Figure 10(a), increasing $\alpha$ initially improves segmentation performance. However, as $\alpha$ continues to increase, performance begins to decline. Empirically, $\alpha=0.5$ achieves a good balance between point and multimodal segmentation performance.